OpenStudy (anonymous):

Calculus: Find the derivative of the following function- -4e^5x/7x+2

OpenStudy (anonymous):

Use the quotient rule.

OpenStudy (anonymous):

\[(\frac{f}{g})'=\frac{gf'-f'g}{g^2}\]

OpenStudy (anonymous):

put \[f(x)=-5e^{5x}\] \[f'(x)=-25e^{5x}\] \[g(x)=7x+2\] \[g'(x)=7\]

OpenStudy (anonymous):

put em in and out pops the answer.

OpenStudy (anonymous):

except of course my formula is wrong!

OpenStudy (anonymous):

Should be:\[(\frac{f}{g})=\frac{f'g-g'f}{g^2}\]

OpenStudy (anonymous):

\[\frac{gf'-fg'}{g^2}\]

OpenStudy (anonymous):

what bendt said! sorry

OpenStudy (anonymous):

What I'm having trouble with is simplifying the answer at the end.

OpenStudy (anonymous):

What do you get?

OpenStudy (anonymous):

\[\frac{-5e^{5x}\times 7 - (7x+2)\times -25e^{5x}}{(7x+2)^2}\]

OpenStudy (anonymous):

numerator is \[-35e^{5x}+175xe^{5x}+40e{5x}\] \[=175xe{5x}+5e^{5x}\] \[=e^{5x}(175x+5)\] \[=5e^{5x}(35x+1)\]unless i made an algebra mistake which is always possible. bendt?

OpenStudy (anonymous):

Can we cheat and use Wolfram Alpha to check?

OpenStudy (anonymous):

ok i made a mistake on the very first line. \[f(x)=-4e^{5x}\] \[f'(x)=-20e^{5x}\]

OpenStudy (anonymous):

i wrote -25 and that was a mistake

OpenStudy (anonymous):

So do you start off by factoring out the e^5x?

OpenStudy (anonymous):

Oh yeah.

OpenStudy (anonymous):

i think my last answer is correct though i will check it since i messed up

OpenStudy (anonymous):

yes it factors out of the whole thing

OpenStudy (anonymous):

yea i screwed up. numerator is \[-4e^{5x}(35x+3)\]

OpenStudy (anonymous):

sorry

OpenStudy (anonymous):

Thank you two very much for your help. This class is very difficult to me :(

OpenStudy (anonymous):

sorry for my error. that -25 threw me off. use -20 and it should be ok

OpenStudy (anonymous):

I'm confused, I think the expression should be:\[\frac{(-20e^{5x})\times(7x+2)-7\times(-4e^{5x})}{(7x-2)^2}\]

OpenStudy (anonymous):

Yes, I think satellite73 figured out the problem using -5e^5x instead of -4e^5x by mistake.

OpenStudy (anonymous):

Ah, ok, as long as you know what your doing now.